A potentiometer consists of a wire of length 4 m and resistance 10W. It is connected to a cell of emf 2V. The potential gradient of wire is


        (a)    0.5 V/m


        (b)    2 V/m


        (c)    5 V/m


        (d)    10 V/m


Grade:10

2 Answers

Aditi Chauhan
9 years ago

(a)

Potential across wire = 2V

Length of wire = 4 m

?
Aman
4 years ago
As we know I max. Is E/R, 
I= E/R  = 2V/10 = 1/5 Ampere
And voltage is I max.  × Resistance ÷ Length ( which is 4m) 
V= I × R/L   =  1×10 / 20 = 0.05V/m